umamsmaiaud:

Ellis is painting wooden fenceposts before putting them in his yard. They are each 7 feet tall and have a diameter of 1 foot. There are 11 fenceposts in all. How much paint will Ellis need to paint all the surfaces of the 11 fenceposts? Use 3.14 for π, and round your answer to the nearest hundredth. Provide an explanation and proof for your answer to receive full credit.

Luigi0210:

Do you know the shape? Otherwise, we could assume that it's a right cylinder

umamsmaiaud:

thats all it said on the question

Luigi0210:

Since they only gave us the diameter and height, we'll assume it's a cylinder.. since it cannot be a cone or rectangle. So for a cylinder, we need to use this formula \[\Large 2\pi*r*h+2\pi*r^2\]

Luigi0210:

You can just plug in values for the surface area formula: \[\large r=0.5 \] \[\large h= 7 \] and \[\large \pi=3.14 \]

umamsmaiaud:

THANK YOUUU

Luigi0210:

Don't forget tho, that will be the surface area of just \(\color{red}{1} \) fence. Since there are 11, just multiply the answer by 11 and you should be set.

Luigi0210:

and of course the units, but you're welcome

umamsmaiaud:

XX

deadz:

:)